# research-article-2col

### v2

@@ 3 Results
The average relaxation times listed in Table 1 for New AMRF measurements can be easily estimated by visual inspection of the relaxation maps in Figure 2.
The center ROI 1 contained a solution of 140 mM NaCl and showed slight variations in uniformity for $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$ and SD maps, but appeared uniform in maps of $\Delta\text{B}_{\text{1}}^{+}$ factor and $\Delta\text{f}_{\text{0}}$. The mean $\text{T}_{\text{1}}$ for New AMRF was 58.9 $\pm$ 1.2 ms versus 57.4 $\pm$ 2.0 ms for RM. The mean $\text{T}_{\text{2,long}}^{*}$ for New AMRF was shorter than $\text{T}_{\text{1}}$ for New AMRF by 18.2 $\pm$ 2.4 ms. For RM, the mean $\text{T}_{\text{2,long}}^{*}$ (53.1 $\pm$ 3.5) was similar to the mean $\text{T}_{\text{1}}$ (57.4 $\pm$ 2.0 ms). The difference between mean $\text{T}_{\text{2,long}}^{*}$ for New AMRF and RM was 11.5 $\pm$ 3.6 ms, where RM was higher. Mean $\text{T}_{\text{2,short}}^{*}$ values for both New AMRF and RM were lower than $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$. The difference between $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,short}}^{*}$ in New AMRF was 26 $\pm$ 2.2 ms while the difference between $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ was 7.7 $\pm$ 2.1 ms. For RM, the difference between $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,short}}^{*}$ was 20.2 $\pm$ 3.4 ms and the difference between $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ was 14.9 $\pm$ 4.4 ms.
@@
SD of the GT was consistently higher than SD from New AMRF. Overlap in the distributions occured only between the 2nd quartile of the GT and 4th quartile of New AMRF in ROI 1, between the 2nd quartile of GT and 3rd quartile of New AMRF in ROI 2, and between the 1st quartile of GT and 4th quartile of New AMRF in ROI 5.
@@
Figure 5 shows the maps for $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$, SD, $\Delta\text{B}_{1}^{+}$ and $\Delta\text{f}_{\text{0}}$ from volunteer 5 in coronal, sagittal and axial slices. We used a weighting factor of k = 20 correlation coefficients during the PC matching process for all volunteers. The most apparent feature in the relaxation times and SD maps was the CSF-filled central ventricles. Contrast from the long $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$ values in CSF dominate the central ventricular structure in all three planes. Contrast from CSF can also be seen within subarachnoid spaces and the cavity along the gyri of the cerebral cortex overall cross sections. In the sagittal slice, we identified the third and fourth ventricles and the occipital horn of the lateral ventricle in maps for $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$ and SD. The overall contrast for $\text{T}_{\text{2,short}}^{*}$ in CSF was reduced compared to $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$. Here the highest values for $\text{T}_{\text{2,short}}^{*}$ were concentrated in the center of the central ventricle in each cross-section. Mean $\text{T}_{\text{2,short}}^{*}$ values reported in Table 3 were the result of this unequal distribution of $\text{T}_{\text{2,short}}^{*}$ over the collective CSF ROI. Mean $\text{T}_{\text{2,short}}^{*}$ in the voxels at the center of the central ventricle across all 5 volunteers was 40.1 $\pm$ 0.1 ms.
@@
The normalized SD maps demonstrated high sodium concentration within CSF and lower sodium concentration in GM and WM. For $\Delta\text{f}_{\text{0}}$, extremely negative and positive shifts were mapped around the medial frontal gyrus, shown in the axial position, over the frontal lobe above the nasal sinus cavity, shown in the sagittal plane, and at the base of the medulla near the posterior cerebellum, shown in the coronal plane.
@@
Table 3 provides a comparison between the relaxation times obtained using our New AMRF method and those reported in the literature at 7 T. For $\text{T}_{\text{1}}$ in CSF measured using our proposed New AMRF, the mean value (59.4 $\pm$ 5.1 ms) was comparable to the mean $\text{T}_{\text{1}}$ reported by Kratzer et al. (19) (61.9 $\pm$ 2.8 ms) using their New AMRF method. $\text{T}_{\text{2}}^{*}$ values for CSF were reported throughout the literature as monoexponential fits with values ranging from 46.3 $\pm$ 6.3 ms (19) to 57.2 $\pm$ 6.6 ms (23). This range indicated that our measured value of 49.7 $\pm$ 6.3 ms for $\text{T}_{\text{2,long}}^{*}$ was in good agreement with the literature. Our mean measurement of 12.5 $\pm$ 3.0 ms for $\text{T}_{\text{2,short}}^{*}$, however, was out of range compared to these same values.
@@
For $\text{T}_{\text{2,long}}^{*}$ in GM, our measured value of 31.5 $\pm$ 2.9 ms was in agreement with previous reports (23, 39, 22, 43), while our measured value of 6.7 $\pm$ 1.0 ms for $\text{T}_{\text{2,short}}^{*}$ was similar to that found by Ridley et al. (39), but slightly longer than other reports (23, 22).
For $\text{T}_{\text{2,long}}^{*}$ in WM, we calculated a mean of 25.2 $\pm$ 3.5 ms, which was within the range of 40.0 $\pm$ 5.2 ms (39) to 22.4 $\pm$ 7.8 ms (23) reported for $\text{T}_{\text{2,long}}^{*}$ in the literature. Our WM $\text{T}_{\text{2,short}}^{*}$ value of 5.5 $\pm$ 0.8 ms was within range of two sources (42, 39) and slightly longer than the other three references (42, 23, 22). Finally, both the $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ values that we measured in GM and WM were in good agreement with values reported for brain tissue (19).
@@
Supplementary Figure S8 shows a series of graphs for the relaxation times measured in each ROI of the 7-compartment phantom plotting the maximum number of matches made to a subset of pixels within a reference range limited by values from the RM. Examining the trends in the graphs for the phantom, we noted that the plot most often changed direction in $\text{T}_{\text{2,long}}^{*}$ from a high number of maximum correlations per pixel match to a global decrease followed by leveling off in the region between k = 10 to k = 50. Direction changes were also noted in $\text{T}_{\text{2,short}}^{*}$ graphs in the neighborhood of k = 100.
4 Discussion
Because $\Delta\text{B}_{\text{1}}^{+}$ and $\Delta\text{f}_{\text{0}}$ were included as dictionary parameters, the spatial influences of transmit inhomogeneity and frequency shift on relaxation times were partially accounted for in the matching process. RF field variations were most pronounced in the phantom data (Figure 2), due to its high average relative permittivity. On the other hand, $\Delta\text{f}_{\text{0}}$ variations were more pronounced in vivo, due to air-filled structures such as the ear canal and maxillary sinus cavities (Figure 6). Some of the artifacts noted in the relaxation maps could be lined up with similar artifacts in maps of $\Delta\text{B}_{\text{1}}^{+}$ and $\Delta\text{f}_{\text{0}}$. This indicated that broadening the ranges for $\Delta\text{B}_{\text{1}}^{+}$ and $\Delta\text{f}_{\text{0}}$ in the dictionary, or reducing the step size for these entries, might improve the appearance of these artifacts in the parameter maps.
@@ 4 Discussion
Changes in the maps of correlation coefficient value averaged over increasing numbers of coefficients were minor. This was because the differences between the maximum value and an average of some subset of coefficients for any single voxel were in the thousandths. Despite such minor changes in value, each correlation represents a potential match between the dictionary and data. We acknowledge that examining correlation in non-convex space and within a noisy environment opens the possibility that some higher valued correlations were calculated for signals representing local minima as opposed to a "true" match. Our choice to include multiple matches based on correlation coeffcient weighting increased the probability that some match resulted from a "true" match as opposed to a local minimum. It may be possible to further improve matching by accounting for non-convexity in the reconstruction process (45, 46, 47).
In the phantom, we weighted the data with 20 coefficients. This reduced the slight inhomogeneity in $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$ and provided the best agreement between New AMRF and the RM data. While weighting beyond 20 coefficients continued to slightly smooth artifacts and improve the agreement between New AMRF and RM in $\text{T}_{\text{1}}$, the agreement to RM in $\text{T}_{\text{2,long}}^{*}$ began to diverge in ROIs 6 and 7 at k = 50. There was no significant change in values beyond the maximum correlation for $\text{T}_{\text{2,short}}^{*}$. Similarly, for SD, an improved overlap in the boxplots between New AMRF and RM is seen at k = 20 with no change beyond this value.
@@
Visually, we could not differentiate between GM and WM in the brain maps of the five volunteers. However, our statistical analysis indicated that the mean $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$ were different enough to distinguish between GM and WM. As listed in Table 3, $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ were reported in several studies where a wide range between the shortest and longest $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ was noted amongst different sources. While different fitting techniques or data acquisition schemes could be the culprit, there may also be variations in relaxation time within GM and WM. Some recent works suggest that regional differences in structural composition within GM and WM (39, 48) are contributing factors, or that local changes are induced within the cellular environment of the respective tissues (3). These considerations, combined with the inherently low SNR of sodium, make distinguishing GM and WM through relaxation mapping difficult.
@@
The ISTO simulation that we used to model $\text{T}_{\text{2}}^{*}$ relaxation included dictionary ranges from 10 to 66 ms for $\text{T}_{\text{2,long}}^{*}$ and from 0.5 to 66 ms for $\text{T}_{\text{2,short}}^{*}$. We hypothesized that average values for $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ in CSF would be equal or very close to each other. However, the mean $\text{T}_{\text{2,short}}^{*}$ was significantly shorter than $\text{T}_{\text{2,long}}^{*}$ in our measurements in CSF. Similar issues with $\text{T}_{\text{2}}^{*}$ fitting in CSF have been mentioned in the literature (49, 23) where the difficulty arose from using discrete biexponential or monoexponential models. Fingerprint matching, however, operates across a pseudo-continuum and is based on the dynamics of the spin system. Theoretically, this would circumvent some of the shortcomings associated with fitting discrete models. The concentration of higher value $\text{T}_{\text{2,short}}^{*}$ in CSF in the center in the central ventricle indicated that partial volume effects from surrounding tissue may have contributed to the low $\text{T}_{\text{2,short}}^{*}$. Eroding the tissue masks did not significantly change the distribution of $\text{T}_{\text{2,short}}^{*}$ in CSF, however, any contamination of the ROI could have resulted in a reduced apparent $\text{T}_{\text{2,short}}^{*}$.
@@
It is also worth noting that our New AMRF method, and the ISTO simulation that we used to generate the fingerprint dictionary, do not make any assumption about either the compartmental homogeneity within the voxel, nor the ratios between relaxation processes for the central and satellite transitions of the New Aspins $\frac{\text{3}}{\text{2}}$ (i.e. the long and short relaxation components). Both New AMRF and the ISTO simulation simply estimate average relaxation times in each voxel, which is very likely a volume weighted average of multiple intra-voxel compartments (extracellular and intracellular spaces from multiple cell types) and local magnetic field inhomogeneities. Consequently, there is very little chance to measure the ideal ratio of 0.6:0.4 for $\text{T}_{\text{2,short}}^{*}$:$\text{T}_{\text{2,long}}^{*}$ due to quadrupolar relaxation processes. Indeed, if we were to specifically measure the ratio of $\text{T}_{\text{2,short}}^{*}$:$\text{T}_{\text{2,long}}^{*}$ with a fitting method (which MRF is not), the ratio of 0.6:0.4 would occur only in a perfectly ideal model where relaxation is purely quadrupolar. This is usually not the case in real samples, including gels, fluids and biological tissues. In reality, dipolar coupling, residual quadrupolar interaction (due to local structural anisotropies), local field inhomogeneities (chemical shift interaction) of various nature, and even spin-orbit interaction where the spin interacts with the magnetic fields generated by the rotational motion of the ion itself, can occur and influence the relaxation process of the New Aspins. Such influences effectively act as pertubations to the main quadrupolar relaxation process, and thus generate New Arelaxation times that deviate from the ideal case of pure quadrupolar relaxation (50, 51, 52).

### v3

@@ 3 Results
Average relaxation times listed in Table 1 for New AMRF measurements can be easily estimated by visual inspection of the relaxation maps in Figure 2. For comparison, RM values are also presented in Table 1 and RM maps are shown in Supplementary Figure S4.
The center ROI 1 contained a solution of 140 mM NaCl and showed slight variations in uniformity for $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$ and SD maps, but appeared uniform in maps of $\Delta\text{B}_{\text{1}}^{+}$ factor and $\Delta\text{f}_{\text{0}}$. The mean $\text{T}_{\text{1}}$ for New AMRF was 58.9 $\pm$ 1.2 ms versus 57.4 $\pm$ 2.0 ms for RM. The mean $\text{T}_{\text{2,long}}^{*}$ for New AMRF was shorter than $\text{T}_{\text{1}}$ for New AMRF by 18.2 $\pm$ 2.4 ms. For RM, the mean $\text{T}_{\text{2,long}}^{*}$ (53.1 $\pm$ 3.5 ms) was similar to the mean $\text{T}_{\text{1}}$ (57.4 $\pm$ 2.0 ms). The difference between mean $\text{T}_{\text{2,long}}^{*}$ for New AMRF and RM was 11.5 $\pm$ 3.6 ms, where RM was higher. Mean $\text{T}_{\text{2,short}}^{*}$ values for both New AMRF and RM were lower than $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$. The difference between $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,short}}^{*}$ in New AMRF was 26 $\pm$ 2.2 ms while the difference between $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ was 7.7 $\pm$ 2.1 ms. For RM, the difference between $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,short}}^{*}$ was 20.2 $\pm$ 3.4 ms and the difference between $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ was 14.9 $\pm$ 4.4 ms.
@@
SD of the GT was consistently higher than SD from New AMRF. Overlap in the distributions occurred only between the 2nd quartile of the GT and 4th quartile of New AMRF in ROI 1, between the 2nd quartile of GT and 3rd quartile of New AMRF in ROI 2, and between the 1st quartile of GT and 4th quartile of New AMRF in ROI 5.
@@
Figure 5 shows the maps for $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$, SD, $\Delta\text{B}_{1}^{+}$ and $\Delta\text{f}_{\text{0}}$ from volunteer 5 in coronal, sagittal and axial slices. We used a weighting factor of k = 20 correlation coefficients during the PC matching process for all volunteers. The most apparent feature in the relaxation times and SD maps was the CSF-filled central ventricles. Contrast from the long $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$ values in CSF dominate the central ventricular structure in all 3 planes. Contrast from CSF can also be seen within subarachnoid spaces and the cavity along the gyri of the cerebral cortex. In the sagittal slice, we identified the third and fourth ventricles and the occipital horn of the lateral ventricle in maps for $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$ and SD. The overall contrast for $\text{T}_{\text{2,short}}^{*}$ in CSF was reduced compared to $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$. Here the highest values for $\text{T}_{\text{2,short}}^{*}$ were concentrated in the center of the central ventricle in each cross-section. Mean $\text{T}_{\text{2,short}}^{*}$ values reported in Table 3 were the result of this unequal distribution of $\text{T}_{\text{2,short}}^{*}$ over the collective CSF ROI. Mean $\text{T}_{\text{2,short}}^{*}$ at the center of the central ventricle across all 5 volunteers was 40.1 $\pm$ 0.1 ms.
@@
The normalized SD maps demonstrated high sodium density within CSF and lower sodium density in GM and WM. For $\Delta\text{f}_{\text{0}}$, extremely negative and positive shifts were mapped around the medial frontal gyrus, shown in the axial position, over the frontal lobe above the nasal sinus cavity, shown in the sagittal plane, and at the base of the medulla near the posterior cerebellum, shown in the coronal plane.
@@
Table 3 provides a comparison between the relaxation times obtained using our New AMRF method and those reported in the literature at 7 T. For $\text{T}_{\text{1}}$ in CSF measured using our New AMRF method, the mean value (59.4 $\pm$ 5.1 ms) was comparable to the mean $\text{T}_{\text{1}}$ reported by Kratzer et al. (20) (61.9 $\pm$ 2.8 ms) using their New AMRF method. Although not included in the Table 3, the $\text{T}_{\text{1}}$ for WM reported using our New AMRF was within the range reported by Coste et al. (33) for WM $\text{T}_{\text{1}}$ at 3 T. $\text{T}_{\text{2}}^{*}$ values for CSF were reported throughout the literature as monoexponential fits with values ranging from 46.3 $\pm$ 6.3 ms (20) to 57.2 $\pm$ 6.6 ms (24). This range indicated that our measured value of 49.7 $\pm$ 6.3 ms for $\text{T}_{\text{2,long}}^{*}$ was in good agreement with the literature. Our mean $\text{T}_{\text{2,short}}^{*}$ of 12.5 $\pm$ 3.0 ms, however, was out of range compared to these same values.
@@
For $\text{T}_{\text{2,long}}^{*}$ in GM, our measured value of 31.5 $\pm$ 2.9 ms was in agreement with previous reports (24, 44, 23, 48), while our measured value of 6.7 $\pm$ 1.0 ms for $\text{T}_{\text{2,short}}^{*}$ was similar to that found by Ridley et al. (44), but slightly longer than other reports (24, 23).
For $\text{T}_{\text{2,long}}^{*}$ in GM, our measured value of 31.5 $\pm$ 2.9 ms was similar to the value determined by Ridley et al. (44) and Fleysher et al. (48) and within the deviation of average values reported by Blunck et al. (24) and Niesporek et al. (23), respectively. Our measured value of 6.7 $\pm$ 1.0 ms for $\text{T}_{\text{2,short}}^{*}$ was similar to that found by Ridley et al. (5.0 $\pm$0.9) (44), but slightly longer than those values reported by Blunck et al. (24) and Niesporek et al. (23). For $\text{T}_{\text{2,long}}^{*}$ in WM, we calculated a mean of 25.2 $\pm$ 3.5 ms, which was within the range of 22.4 $\pm$ 7.8 ms (24) to 40.0 $\pm$ 5.2 ms (44) reported for $\text{T}_{\text{2,long}}^{*}$ in the literature and provided in Table 3. In this case, our value was similar to those reported by Blunck et al. (24) and Niesporek et al. (23), but lower than those reported by Lommen et al.(47), Ridley et al. (44) and Fleysher et al. (48), respectively. Our WM $\text{T}_{\text{2,short}}^{*}$ value of 5.5 $\pm$ 0.8 ms our values were similar to that reported by Lommen et al. (47), within the deviation of the reported value from Ridley et al.(44) but higher than those reported by Blunck et al. (24) and Niesporek et al.(23), respectively. Finally, both the $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ values that we measured in GM and WM were in good agreement with values reported for brain tissue by Kratzer et al. (20), which the authors provided as GM and WM combined.
@@
Supplementary Figure S9 shows a series of graphs for the relaxation times measured in each ROI of the 7-compartment phantom plotting the maximum number of matches made to a subset of pixels within a reference range limited by values from the RM. Examining the trends in the graphs, we noted that the plot most often changed direction in $\text{T}_{\text{2,long}}^{*}$ from a high number of maximum correlations per pixel match to a global decrease followed by leveling off in the region between k = 10 to k = 50. Direction changes were also noted in $\text{T}_{\text{2,short}}^{*}$ graphs in the neighborhood of k = 100.
4 Discussion
Because $\Delta\text{B}_{\text{1}}^{+}$ and $\Delta\text{f}_{\text{0}}$ were included as dictionary parameters, the spatial influences of transmit inhomogeneity and frequency shift on relaxation times were partially accounted for in the matching process. RF field variations were most pronounced in the phantom data (Figure 2), due to its high average relative permittivity. On the other hand, $\Delta\text{f}_{\text{0}}$ variations were more pronounced in vivo, due to air-filled structures such as the ear canal and maxillary sinus cavities (Figure 6). Some of the artifacts noted in the relaxation maps could be lined up with similar artifacts in maps of $\Delta\text{B}_{\text{1}}^{+}$ and $\Delta\text{f}_{\text{0}}$. This alignment of artifacts demonstrated how the spatial influence of transmit inhomogeneity and frequency shift were articulated in the relaxation maps through the matching process, and how inhomogeneity influenced the matching of signals in the relaxation maps. These results indicated that broadening the ranges for $\Delta\text{B}_{\text{1}}^{+}$ and $\Delta\text{f}_{\text{0}}$ in the dictionary, or reducing the step size for these entries, might improve the appearance of these artifacts in the parameter maps. For this study, our choice of $\Delta\text{B}_{\text{1}}^{+}$ factor and $\Delta\text{f}_{\text{0}}$ steps reflect a compromise between computation time and the possible range for $\Delta\text{B}_{\text{1}}^{+}$ factor and $\Delta\text{f}_{\text{0}}$ variation in head at 7T. Also significant is that our New AMRF method does not use the maps for $\Delta\text{B}_{\text{1}}^{+}$ and $\Delta\text{f}_{\text{0}}$ generated through the matching process as a post-processing correction to the final relaxation maps. All maps produced using our New AMRF technique ($\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$, $\text{T}_{\text{2,short}}^{*}$ and sodium density) were generated from matching the dictionary of simulated signals to real data where $\Delta\text{B}_{1}^{+}$ and $\Delta\text{f}_{\text{0}}$ corrections were built into the dictionary.
@@ 4 Discussion
Changes in the maps of correlation coefficient value averaged over increasing numbers of coefficients were minor. This was because the differences between the maximum value and an average of some subset of coefficients for any single voxel were in the thousandths. Despite such minor changes in value, each correlation represents a potential match between the dictionary and data. We acknowledge that examining correlation in non-convex space and within a noisy environment opens the possibility that some higher valued correlations were calculated for signals representing local minima as opposed to a "true" match. Our choice to include multiple matches based on correlation coeffcient weighting increased the probability that some match resulted from a "true" match as opposed to a local minimum. It may be possible to further improve matching by accounting for non-convexity in the reconstruction process (50, 51, 52).
[…]
Moreover, while results in phantoms show quite good agreement between our proposed New AMRF method and RM, we acknowledge the difficulty of obtaining accurate relaxation times with New AMRI in general, as most values are within 5 ms here, which is likely within SNR variations between the methods that we compared, since New Aimaging is inherently a low-SNR technique, even for RM.
@@
Previous studies have indicated that differences in tissue structure and mobility affect the residual quadrupolar interaction arising from New Aspins interacting with their environment(2). More restricted environments, such as myelinated white matter, are associated with stronger residual quadrupolar couplings that lead to faster decay components. However, there is also evidence that sodium exists as different pools of nuclei within individual brain tissues, where some nuclei exhibit biexponential relaxation while other spin pool distributions demonstrate monoexponential behavior(26). Other works support similar considerations that regional differences in structural composition within GM and WM (44, 53), or local changes induced within the cellular environment of the GM and WM (3) confound the ability to explicitly distinguish between these tissues, respectively. Combined with the inherently low SNR of sodium imaging, these factors make distinguishing GM and WM through relaxation mapping difficult. In evaluating the ability of our New AMRF method to distinguish between GM and WM we compared our results against 7 references (Table 3) where $\text{T}_{\text{1}}$, $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ were reported as mean values across a cohort of volunteers in each study. Using similar mean value analysis, we reported comparable results in the last line of Table 3, which indicated that our method could distinguish between GM and WM as well as other methods listed. Additionally, we were able to demonstrate that the mean $\text{T}_{\text{1}}$ and $\text{T}_{\text{2,long}}^{*}$ were different enough to distinguish between GM and WM using statistical analysis, provided in the Results section. Despite the statistical analysis and mean value support for previous findings, visual inspection of the brain maps of the five volunteers alone was not sufficient for differentiating between GM and WM.
For those references considered in Table 3 where $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ were reported, there is a wide range between the shortest and longest $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ noted between different sources. Different fitting techniques or data acquisition schemes could be the culprit as well as some variation in relaxation time within GM and WM.
Examining the references in Table 3, we noted that several acquisition schemes were in play by different authors; Blunck et al. (24)., Lommen et al. (47), Ridley et al. (44), Niesporek et al. (23) and Nagel et al. (22) applied variations of multiple echo acquisitions with varied TE and 3D radial readouts. Fleysher et al. (48) used a gradient recalled echo (GRE) sequence with a Cartesian readout. In these works, the authors determined $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ using using voxel wise fitting in image space where a signal model was defined as biexponential decay and the signal contributions from satellite and central transitions were expected to demonstrate a $\text{T}_{\text{2,short}}^{*}$:$\text{T}_{\text{2,long}}^{*}$ ratio of 0.6:0.4. In some work, this ratio was explicitly built-in to the fitting model itself(24, 47, 23, 48). Kratzer and colleagues have previously reported two methods for New AMRF. The first involved using modified Bloch equations as a simplified simulation for the New Aspin $\frac{3}{2}$ system, was capable of single slice 2D acquisition and suffered from low efficiency (19). Their followup to this work utilized an ISTO simulation technique and was expanded to 3D (20). In our study we also employed the ISTO formalism to model New Arelaxation. Our New AMRF presented here is an extension of the methods applied in our previous work (21). In order to capture the $\text{T}_{\text{2,short}}^{*}$ dynamics of the New Aspin system, several of the works compared in Table 3 utilized a 3D radial trajectory (24, 47, 23, 22), including Kratzer et al. (20) who used a center-out radial trajectory with variable TE, incorporating both single and double echo readouts. In our sequence we utilized the 3D FLORET trajectory (38, 25, 29, 21). FLORET has several advantages over radial acquisition schemes, including improved image quality due to fast gradient spoiling allowing for shorter minimal TR, and improved SNR due to more uniform sampling (29). The goal of our pulse sequence optimization in developing the New AMRF train (Figure 1) was to best differentiate between GM and WM. This was accomplished by minimizing the PC between the simulated tissue systems using a GA. Our approach resulted in a significantly shorter pulse train than that reported by Kratzer et al. (19, 20), who optimized FAs using the Cramer-Rao lower bound (CRLB) criterion in development of a 1000-pulse train. Our dictionary was generated through direct numerical simulation of the density matrix dynamics using ISTO and incorporated explicit modeling of a $\Delta\text{B}_{1}^{+}$ factor and $\Delta\text{f}_{\text{0}}$ shift to account for the influence of inhomogeneities directly into the system as described in the Methods section. Kratzer et. al. also used ISTO formalism in their dictionary simulation and included a $\Delta\text{B}_{0}$ correction interpolated along the $\bar{\Delta\text{B}_{0}}$ axis. These authors extracted relaxation parameters through matching using the highest scalar product between the compressed dictionary and reconstructed images (20). Our New AMRF matching differs significantly. First, we matched an uncompressed dictionary to reconstructed image data using the PC. Second, we employed a correlation-weighted strategy to include multiple high correlation dictionary matches per voxel.
In our previous work we demonstrated that the evolution of the ISTOs represent the full spin dynamics of the New Aspin $\frac{3}{2}$ system, where the simulation of system evolution and relaxation accounts for all single, double and triple quantum coherences (SQC, DQC, TQC). Although multiquantum coherence (MQC) can evolve during a multipulse sequence, and affect the final magnetization (21), the design of our New AMRF pulse sequence did not specifically filter for MQCs. Other methods have been developed for selective filtering of both DQCs (2, 54) and TQCs (2). In each case, careful phase cycling and RF pulse FA selection were designed to select for either DQC or TQC (2). In contrast to this approach, our New AMRF pulse train design was motivated for separating GM and WM signals based on relaxation properties and was developed through combined ISTO simulation of the New Aensemble with GA optimization.
@@
In our New AMRF method, the ISTO simulation that we used to model $\text{T}_{\text{2}}^{*}$ relaxation included dictionary ranges from 10 to 66 ms for $\text{T}_{\text{2,long}}^{*}$ and from 0.5 to 66 ms for $\text{T}_{\text{2,short}}^{*}$. We hypothesized that average values for $\text{T}_{\text{2,long}}^{*}$ and $\text{T}_{\text{2,short}}^{*}$ in CSF would be equal or very close to each other. However, the mean $\text{T}_{\text{2,short}}^{*}$ was significantly shorter than $\text{T}_{\text{2,long}}^{*}$ in our measurements in CSF. Similar issues with $\text{T}_{\text{2}}^{*}$ fitting in CSF have been mentioned in the literature (55, 24) where the difficulty arose from using discrete biexponential or monoexponential models. Fingerprint matching, however, operates across a pseudo-continuum and is based on the dynamics of the spin system. Theoretically, this would circumvent some of the shortcomings associated with fitting discrete models. The concentration of higher value $\text{T}_{\text{2,short}}^{*}$ in CSF in the center in the central ventricle indicated that partial volume effects from surrounding tissue may have contributed to the low $\text{T}_{\text{2,short}}^{*}$. Eroding the tissue masks did not significantly change the distribution of $\text{T}_{\text{2,short}}^{*}$ in CSF, however, any contamination of the ROI could have resulted in a reduced apparent $\text{T}_{\text{2,short}}^{*}$.
@@
It is also worth noting that our New AMRF method, and the ISTO simulation that we used to generate the fingerprint dictionary, do not make any assumption about either the compartmental homogeneity within the voxel, nor the ratios between relaxation processes for the central and satellite transitions of the New Aspins $\frac{\text{3}}{\text{2}}$ (i.e. the long and short relaxation components). Both New AMRF and the ISTO simulation simply estimate average relaxation times in each voxel, which is very likely a volume weighted average of multiple intra-voxel compartments (extracellular and intracellular spaces from multiple cell types) and local magnetic field inhomogeneities. Consequently, there is very little chance to measure the ideal ratio of 0.6:0.4 for $\text{T}_{\text{2,short}}^{*}$:$\text{T}_{\text{2,long}}^{*}$ due to quadrupolar relaxation processes. Indeed, if we were to specifically measure the ratio of $\text{T}_{\text{2,short}}^{*}$:$\text{T}_{\text{2,long}}^{*}$ with a fitting method (which MRF is not), the ratio of 0.6:0.4 would occur only in a perfectly ideal model where relaxation is purely quadrupolar. This is usually not the case in real samples, including gels, fluids and biological tissues. In reality, dipolar coupling, residual quadrupolar interaction (due to local structural anisotropies), local field inhomogeneities (chemical shift interaction) of various nature, and even spin-orbit interaction where the spin interacts with the magnetic fields generated by the rotational motion of the ion itself, can occur and influence the relaxation process of the New Aspins. Such influences effectively act as pertubations to the main quadrupolar relaxation process, and thus generate New Arelaxation times that deviate from the ideal case of pure quadrupolar relaxation (56, 35, 57).